# Conformally invariant boundary arcs in double dimers

### v1

@@ 1. Introduction
Our motivation (outside of the dimer model) comes from the theory of the continuum Gaussian free field (GFF) and its level sets. The GFF with Dirichlet boundary conditions in a domain $D$ is a Gaussian field whose covariance kernel is given by the Green’s function $g(x,y)$ of Brownian motion in the domain with Dirichlet boundary condition. In this article we assume that the Green’s function is normalised so that $g(x,y)\sim(-2\pi)^{-1}\log|x-y|$ as $x\to y$ in the interior of the domain. We are guided by the construction of Qian and Werner [19] who coupled two different versions of the GFF, with Dirichlet and Neumann (free) boundary conditions (we refer the reader to [19] for a definition of the latter), through a common set of level lines that form an ALE.
Even though the GFF is not a function, but only a generalised function (a Schwartz distribution acting on test functions), a beautiful theory of its level sets has been developed in recent years [4, 3, 2]. An important notion in this geometric representation of the GFF is that of the two-valued sets $\mathbb{A}_{-a,b}$, where $a,b>0$, $a+b\geq 2\lambda$, and where $\lambda=\sqrt{{\pi}/8}$ is the constant appearing in the height-gap phenomenon of Schramm and Sheffield [21]. Formally they are thin local sets in the sense of [4] whose associated harmonic function takes only two values $-a$ and $b$. Heuristically they are random sets naturally coupled with the GFF that could be seen as sets of those points in $D$ that are connected to $\partial D$ by a path on which the values of the field (which is not defined pointwise) lie between $-a$ and $b$. The ALE is then the set of arcs given by the (restriction to $D$ of the) connected components of $A_{-\lambda,\lambda}$. In [19] a surprising coupling between the ALE and two different types of the GFF was given. In this construction, the Dirichlet GFF (with zero boundary conditions) changes its value by ±$\pm 2\lambda$ across the arcs of the ALE in a deterministic and alternating fashion, whereas the Neumann (free boundary condition) GFF changes its value by ±$\pm 2\lambda$ in an i.i.d. fashion.
@@
Figure 1. Left: a (partial) dimer cover of $\mathcal{G}_{\varepsilon}^{r}$ is represented in red, and the lower part is folded on the upper part in dotted red. A folded dimer configuration appears on the upper part: loops and arcs alternate between full red and dotted red. The graph $\mathcal{G}_{\varepsilon}^{r}$ has Temperleyan boundary conditions: all its corners are black squares, and one of the black squares on the boundary (in this case on the right-hand side of the intersection with the real line) is removed. Right: (partial) dimer covers of $\mathcal{G}_{\varepsilon}^{r}\cap\mathbb{R}\times\mathbb{R}_{\geq 0}$ and $\mathcal{G}_{\varepsilon}^{r}\cap\mathbb{R}\times\mathbb{R}_{<0}$ are represented in solid red. The lower part is superimposed on the upper part in dotted red. A superimposed configuration appears on the upper part: loops and arcs alternate between full red and dotted red. The upper part has again Temperleyan boundary conditions, and the lower part has piecewise Temperleyan boundary conditions with two white bullet corners.
@@
In this article we consider a set of discrete arcs in two different but related versions of double dimers that satisfy an analogous coupling when considered together with the associated height function (that is known to converge to the Gaussian free field in a variety of settings [16, 20, 6, 14]). To be precise, recall that the dimer model is a random perfect matching (which in the setting of this work is chosen uniformly at random) of a finite graph. We will assume that the graph is a finite connected subgraph of the rescaled square lattice $\epsilon\mathbb{Z}^{2}$ that is symmetric across the real line, and whose restriction to the upper half-plane approximates a simply connected continuum domain $U$. We will denote such graphs by $\mathcal{G}_{\varepsilon}^{r}$.
@@
In the first setup, referred to as folded dimers, that corresponds to the Neumann GFF in the coupling of [19], we consider the dimer model on $\mathcal{G}_{\varepsilon}^{r}$, and fold the resulting perfect matching along the real line as in Fig. 1. This results in a collection of loops, doubled edges, and arcs starting and ending on the real line. The associated height function on the faces of the folded graph can be defined (up to a global additive constant) by assigning orientations to each dimer in the folded dimer cover: each dimer coming from the top half of $\mathcal{G}_{\varepsilon}^{r}$ is oriented from a black to a white vertex (in a fixed bipartite colouring of the lattice), and each dimer coming from the bottom half is oriented from a white vertex to a black vertex. This results in a consistent orientation of each loop and arc, and finally one declares that the increment of the height function from a face $u$ to a neighbouring face is $+1$ (resp. $-1$) if there is a loop or an arc separating the two faces such that $u$ lies on the left-hand (resp. right-hand) side of the oriented dimer in the loop or arc. Note each loop and arc can be oriented in the two possible ways with equal probability and independently of other loops and arcs. This corresponds to the increment of the height function being ±$\pm 1$, and is analogous to the coupling of Qian and Werner between the ALE and the GFF with mixed Neuman/Dirichlet boundary conditions.
@@
One can see that the height function in this setting does indeed converge to the continuum GFF with mixed Neuman/Dirichlet boundary conditions. Indeed, it is known from the original work of Kenyon [16] (and more recent generalisations [20, 6]) that if the dimer model on $\mathcal{G}_{\varepsilon}^{r}$ has Temperleyan boundary conditions (which we assume to be the case in our arguments as is shown in Fig. 1 and Fig. 11), then the centered height function of the dimer model on $\mathcal{G}_{\varepsilon}^{r}$, multiplied by $2\lambda$ to match the height gap across the continuum level lines, converges as $\varepsilon\to 0$ to $1/\sqrt{2}$ times the Dirichlet GFF in the symmetric domain $U^{r}$ whose restriction to the upper half plane is $U$ (i.e. $U^{r}=U\cup\bar{U}\cup(\partial U\cap\{0\}\times\mathbb{R})$). We note that here we take into account the fact that the height function of a single dimer model defined in [16] is $4$ times the height function we are using in this article. From this definition it is clear that the expectation of the height function on $\mathcal{G}_{\varepsilon}^{r}$ is antisymmetric across the real line, and that the height function at a face $u$ in the folded dimer model is by definition equal to the sum of the height functions at $u$ and at its reflection $\bar{u}$ in the single dimer model on $\mathcal{G}_{\varepsilon}^{r}$. Since the continuum GFF in $U$ with Neumann boundary conditions on the real line and Dirichlet boundary conditions on the rest of $\partial U$ can be obtained in the same way from $1/\sqrt{2}$ times the Dirichlet GFF in $U^{r}$ (as e.g. described in [19, 7]), we can conclude that $2\lambda$ times the height function defined by the discrete loops and arcs converges to exactly the same GFF with mixed boundary conditions as in the coupling of [19].
@@
In the second setup, we first remove the row of vertical edges of $\mathcal{G}_{\varepsilon}^{r}$ just below the real line (see Fig. 1) and then repeat the folding procedure as in the first setup. This results in a number of differences compared to the previous case. The first one is that now the arcs emanating from the real line cannot have arbitrary orientations. Indeed, one quickly realises that due to parity reasons the arcs must deterministically alternate in orientation – just as in the part of the coupling in [19] concerning the Dirichlet GFF. Moreover, after the removal of edges the graph splits into two connected components $\mathcal{G}_{\varepsilon}^{r}\cap\mathbb{R}\times\mathbb{R}_{\geq 0}$ and $\mathcal{G}_{\varepsilon}^{r}\cap\mathbb{R}\times\mathbb{R}_{<0}$. One can check that the the upper component has again Temperleyan boundary conditions, whereas the lower one has piecewise Temperleyan boundary conditions (see Fig. 1) introduced by Russkikh [20], who proved convergence of $2\lambda$ times the centered height function to $1/\sqrt{2}$ times the Dirichlet GFF. Later Berestycki and Liu showed [8, Theorem 5.1] that the mean of $2\lambda$ times the height function converges to a harmonic function with piecewise constant boundary conditions with alternating jumps of size ±$\pm\lambda$ located at the points of the Temperleyan corners (in our case these are the two extremities of $\partial U\cap\{0\}\times\mathbb{R}$) plus a term that involves the winding of the boundary of $U$. Since $U$ and $\bar{U}$ have opposite windings of the boundary, they cancel out when the two height functions are added, and we can conclude that the mean of $2\lambda$ times the height function of the superimposed configuration converges to the harmonic function in $U$ with $\lambda$ boundary conditions on $\partial U\cap\{0\}\times\mathbb{R}$ and zero elsewhere. Since the sum of two independent fields that are both $1/\sqrt{2}$ times the Dirichlet GFF (coming from the upper and lower domains) is ($1$ times) the Dirichlet GFF, we can conclude that $2\lambda$ times the height function of our model converges to the corresponding field described in [19], i.e. a GFF in $U$ with boundary conditions equal to $\lambda$ on $\partial U\cap\{0\}\times\mathbb{R}$ and zero on the rest of $\partial U$.
@@
We conjecture (on top of the conjecture of [7] that concerns the first model) that the arcs in both models converge to the ALE in $U$ (emanating from $\partial U\cap\{0\}\times\mathbb{R}$ as in [19]). This is strongly evidenced by the discussion above. However, convergence of the height function is known not to be enough to conclude convergence of interfaces in the double dimer model [17, 13, 5]. Finally let us mention that the two models yield different measures in the discrete setting, when one forgets the orientation of loops and arcs. Indeed, in the first model each nontrivial arc (that is not a single edge) comes with a combinatorial factor of two corresponding to the two different orientations, and it comes with a factor of one in the second model. The number of nontrivial arcs is not constant (unlike the number of all arcs) and hence the measures are different. Nonetheless we conjecture that the two collections of arcs have the same scaling limit.
@@
The main observables of interest in both models will be $n_{\varepsilon}(z)$ – the number of arcs that surround a point $z$, i.e. arcs that separate $z$ from the part of the boundary with no arcs $\partial U\setminus\{0\}\times\mathbb{R}$, and $o_{\varepsilon}(z)$ – the parity of $n_{\varepsilon}(z)$. Our main result can be summarised as follows.
@@
One of the main conceptual difficulties in approaching the problem of studying the arcs themselves is the issue of separating the loops from the arcs in the counting arguments. This is achieved by developing a novel Kenyon’s formula for double dimers with arcs, and afterwards choosing the right weights (or more precisely SL(2) connections) in this formula. Afterwards an involved analysis of the scaling limit of the formula by means of discrete complex analysis techniques is required.
We note that the works [17, 13, 5] study multi-point functions of topological observables in the double dimer model. Basok and Chelak [5] showed that these observables contain enough information to identify the law of the loops. In our case, the fact that we want to disregard all the loops from the picture restricts greatly the number of useful topological observables, and we can only treat the one-point (or two-point) function. However, it will be very interesting to develop the methods of [17, 13, 5] to study the full collection of both arcs and loops, that (based on the conjecture of Kenyon [17]) and our conjectures should converge to the ALE with an independent CLE4 in its complement.
@@ Acknowledgements
The authors extend their warm thanks to Avelio Sepúlveda for his contribution of Appendix A. ML and LR thank Misha Basok for useful discussions during the program Geometry, Statistical Mechanics, and Integrability at IPAM, UCLA. ML and KR were supported by the FWF Standalone grant Spins, Loops and Fields P 36298 and the SFB Discrete Random Structures F 1002. KR was supported by the Academy of Finland Centre of Excellence Programme grant number 346315 “Finnish centre of excellence in Randomness and STructures” (FiRST). LR was partially supported by the DIMERS project ANR-18-CE40-0033 funded by the French National Research Agency.
@@ 2.1. Setup and formula
We begin with some definitions. Let $G=(V,E)$ be a finite, connected, bipartite, planar graph. Its vertices $V$ are partitioned into white and black vertices. Let $\partial$ be a strict subset of the vertices incident to the external face of $G$, which contains the same number of black and white vertices. We form a graph $G^{\times}$ which can be described as taking two copies of $G$, identifying the vertices in the two copies on $\partial$, and then adding “diagonal” edges which join a vertex $u$ in one copy of $G$ to a vertex $v$ in the other copy, whenever $u$ and $v$ form an edge in $G$. One can think of the two copies of $G$ being “folded” together along $\partial$.
@@
[…]
Let $\Phi$ be some $\mathrm{SL}_{2}(\mathbb{C})$ connection on the bulk of $G$. That is, $\Phi$ assigns each directed edge $uv$ ($u,v\in V\setminus\partial$) a matrix $\phi_{u,v}\in\mathrm{SL}_{2}(\mathbb{C})$, such that $\phi_{v,u}=\phi_{u,v}^{-1}$. For a loop $C=\{e_{1},\dots,e_{2n}\}$ in $G$, define the monodromy of $\Phi$ around $C$, $\phi_{C}$ to be the product $\phi_{e_{1}}\circ\cdots\circ\phi_{e_{2n}}$. Note this depends on the choice of $e_{1}$ and the orientation or the loop; $\mathrm{tr}(\phi_{C})$, however, does not. For an edge $\{u,w\}$ with $w\in\partial$ and $u\in V\setminus\partial$, let $\psi_{\{u,w\}}\in\mathbb{C}^{2}$. Fix some basis of each copy of $\mathbb{C}^{2}$ so that each $\phi_{v,u}$ can be written as a matrix, and each $\psi_{\{u,w\}}$ as a column vector.
@@
It is well known that such phases exist (see for example Theorem 4.1 of [10]) and in fact can be taken to be real i.e. signs ±$\pm 1$ on the edges. One can show that (1) holds for all loops surrounding an even number of vertices. Moreover if one replaces $\xi_{x,y}$ by $\varepsilon(x)\varepsilon(y)\xi_{x,y}$ for some function $\varepsilon:V\to\{\pm 1\}$, the new phases are also a set of Kasteleyn phases: this is called a gauge transform. Let $\bar{\partial}$ be a connected, strict subset of the boundary vertices, with $\partial\subset\bar{\partial}$. By using a guage transform, one can assume that the phases on the boundary edges (incident to the external face) alternate between $+1$ and $-1$ in $\bar{\partial}$. We fix such real phases $\xi$ for the rest of this section.
@@
We assign an ordering to $V^{\times}$ which will be used in the Pfaffian of Proposition 2.1. Let the ordering be: first the boundary vertices, ordered as they appear clockwise in $\overline{\partial}$, then the white vertices in some order such that each $(u,1)$ is directly followed by $(u,2)$, and then the black vertices, again in some order such that $(v,1)$ is directly followed by $(v,2)$. We can write our assumption that the Kasteleyn phases alternate on the boundary as $\xi_{w,b}=(-1)^{\mathbbm{1}\{w>b\}}$ for all edges linking two vertices of $\overline{\partial}$.
@@
Figure 2. An example of a graph $G$, where we take $\partial=\overline{\partial}$ to be the edges on the $x$-axis, and real Kasteleyn weights $\xi$ which alternate in sign along $\overline{\partial}$ as $\xi_{w,b}=(-1)^{\mathbbm{1}\{w>b\}}$, where the ordering on $\partial$ increases clockwise. (We only write the $-1$ phases in the diagram for clarity; the rest are 1).
@@
(thinking of $b,w\in\mathbb{C}$), that is, at each white vertex, the weights of the four adjacent edges starting at the right-going edge and proceeding anticlockwise are $1,i,-1,-i$ respectively. Using the gauge function $\varepsilon(u)=(-i)^{\mathbbm{1}\{u\text{ has odd vertical coordinate}\}}(-1)^{\mathbbm{1}\{u\ \text{is white}\}}$, and the real Kasteleyn phases $\xi$ depicted in Figure 2, $\zeta_{u,v}:=\varepsilon(u)\varepsilon(v)\xi_{u,v}$ are the Kasteleyn phases of discrete holomorphy of Equation (3). Remark 2.3 gives us the corollary of Proposition 2.1 which we will use in the following sections.

### v2

@@ 1. Introduction
Our motivation (outside of the dimer model) comes from the theory of the continuum Gaussian free field (GFF) and its level sets. The GFF with Dirichlet boundary conditions in a domain $D$ is a Gaussian field whose covariance kernel is given by the Green function $g(x,y)$ of Brownian motion in the domain with Dirichlet boundary condition. In this article we assume that the Green function is normalised so that $g(x,y)\sim(-2\pi)^{-1}\log|x-y|$ as $x\to y$ in the interior of the domain. We are guided by the construction of Qian and Werner [18] who coupled two different versions of the GFF, with Dirichlet and Neumann (free) boundary conditions (we refer the reader to [18] for a definition of the latter), through a common set of level lines that form an ALE.
Even though the GFF is not a function, but only a generalised function (a Schwartz distribution acting on test functions), a beautiful theory of its level sets has been developed in recent years [4, 3, 2]. An important notion in this geometric representation of the GFF is that of the two-valued sets $\mathbb{A}_{-a,b}$, where $a,b>0$, $a+b\geq 2\lambda$, and where $\lambda=\sqrt{{\pi}/8}$ is the constant appearing in the height-gap phenomenon of Schramm and Sheffield [20]. Formally they are thin local sets in the sense of [4] whose associated harmonic function takes only two values a-a and $b$. Heuristically they are random sets naturally coupled with the GFF that could be seen as sets of those points in $D$ that are connected to $\partial D$ by a path on which the values of the field (which is not defined pointwise) lie between a-a and $b$. The ALE is then the set of arcs given by the (restriction to $D$ of the) connected components of $A_{-\lambda,\lambda}$. In [18] a surprising coupling between the ALE and two different types of the GFF was given. In this construction, the Dirichlet GFF (with zero boundary conditions) changes its value by $\pm 2\lambda$ across the arcs of the ALE in a deterministic and alternating fashion, whereas the Neumann (free boundary condition) GFF changes its value by $\pm 2\lambda$ in an i.i.d. fashion.
@@
Figure 1. Left: a (partial) dimer cover of $\mathcal{G}_{\varepsilon}^{r}$ is represented in red, and the lower part is folded on the upper part in dotted red. A folded dimer configuration appears on the upper part: loops and arcs alternate between full red and dotted red. The graph $\mathcal{G}_{\varepsilon}^{r}$ has Temperleyan boundary conditions: all its corners are black squares, and one of the black squares on the boundary (in this case on the right-hand side of the intersection with the real line) is removed. Right: (partial) dimer covers of $\mathcal{G}_{\varepsilon}^{r}\cap(\mathbb{R}\times\mathbb{R}_{\geq 0})$ and $\mathcal{G}_{\varepsilon}^{r}\cap(\mathbb{R}\times\mathbb{R}_{<0})$ are represented in solid red. The lower part is superimposed on the upper part in dotted red. A superimposed configuration appears on the upper part: loops and arcs alternate between full red and dotted red. The upper part has again Temperleyan boundary conditions, and the lower part has piecewise Temperleyan boundary conditions with two white bullet corners.
@@
In this article we consider a set of discrete arcs in two different but related versions of double dimers that satisfy an analogous coupling when considered together with the associated height function (that is known to converge to the Gaussian free field in a variety of settings [15, 19, 6, 13]). To be precise, recall that the dimer model is a random perfect matching (which in the setting of this work is chosen uniformly at random) of a finite graph. We will assume that the graph is a finite connected subgraph of the rescaled square lattice $\varepsilon\mathbb{Z}^{2}$ that is symmetric across the real line, and whose restriction to the upper half-plane approximates a simply connected continuum domain $U$. We will denote such graphs by $\mathcal{G}_{\varepsilon}^{r}$.
@@
In the first setup, referred to as folded dimers, that corresponds to the Neumann GFF in the coupling of [18], we consider the dimer model on $\mathcal{G}_{\varepsilon}^{r}$, and fold the resulting perfect matching along the real line as in Fig. 1. This results in a collection of loops, doubled edges, and arcs starting and ending on the real line. The associated height function on the faces of the folded graph can be defined (up to a global additive constant) by assigning orientations to each dimer in the folded dimer cover: each dimer coming from the top half of $\mathcal{G}_{\varepsilon}^{r}$ is oriented from a black to a white vertex (in a fixed bipartite colouring of the lattice), and each dimer coming from the bottom half is oriented from a white vertex to a black vertex. This results in a consistent orientation of each loop and arc, and finally one declares that the increment of the height function from a face $u$ to a neighbouring face is +1+1 (resp. 1-1) if there is a loop or an arc separating the two faces such that $u$ lies on the left-hand (resp. right-hand) side of the oriented dimer in the loop or arc. Note each loop and arc can be oriented in the two possible ways with equal probability and independently of other loops and arcs. This corresponds to the increment of the height function being $\pm 1$, and is analogous to the coupling of Qian and Werner between the ALE and the GFF with mixed Neuman/Dirichlet boundary conditions.
@@
One can see that the height function in this setting does indeed converge to the continuum GFF with mixed Neuman/Dirichlet boundary conditions. Indeed, it is known from the original work of Kenyon [15] (and more recent generalisations [19, 6]) that if the dimer model on $\mathcal{G}_{\varepsilon}^{r}$ has Temperleyan boundary conditions (which we assume to be the case in our arguments as is shown in Fig. 1 and Fig. 12), then the centered height function of the dimer model on $\mathcal{G}_{\varepsilon}^{r}$, multiplied by $2\lambda$ to match the height gap across the continuum level lines, converges as $\varepsilon\to 0$ to $1/\sqrt{2}$ times the Dirichlet GFF in the symmetric domain $U^{r}$ whose restriction to the upper half plane is $U$ (i.e. $U^{r}=U\cup\bar{U}\cup(\partial U\cap(\{0\}\times\mathbb{R}))$). We note that here we take into account the fact that the height function of a single dimer model defined in [15] is $4$ times the height function we are using in this article. From this definition it is clear that the expectation of the height function on $\mathcal{G}_{\varepsilon}^{r}$ is antisymmetric across the real line, and that the height function at a face $u$ in the folded dimer model is by definition equal to the sum of the height functions at $u$ and at its reflection $\bar{u}$ in the single dimer model on $\mathcal{G}_{\varepsilon}^{r}$. Since the continuum GFF in $U$ with Neumann boundary conditions on the real line and Dirichlet boundary conditions on the rest of $\partial U$ can be obtained in the same way from $1/\sqrt{2}$ times the Dirichlet GFF in $U^{r}$ (as e.g. described in [18, 7]), we can conclude that $2\lambda$ times the height function defined by the discrete loops and arcs converges to exactly the same GFF with mixed boundary conditions as in the coupling of [18].
@@
In the second setup, which we refer to as shifted dimers, we first remove the row of vertical edges of $\mathcal{G}_{\varepsilon}^{r}$ just below the real line (see Fig. 1) and then repeat the folding procedure as in the first setup. This results in a number of differences compared to the previous case. The first one is that now the arcs emanating from the real line cannot have arbitrary orientations. Indeed, one quickly realises that due to parity reasons the arcs must deterministically alternate in orientation – just as in the part of the coupling in [18] concerning the Dirichlet GFF. Moreover, after the removal of edges the graph splits into two connected components $\mathcal{G}_{\varepsilon}^{r}\cap(\mathbb{R}\times\mathbb{R}_{\geq 0})$ and $\mathcal{G}_{\varepsilon}^{r}\cap(\mathbb{R}\times\mathbb{R}_{<0})$. One can check that the the upper component has again Temperleyan boundary conditions, whereas the lower one has piecewise Temperleyan boundary conditions (see Fig. 1) introduced by Russkikh [19], who proved convergence of $2\lambda$ times the centered height function to $1/\sqrt{2}$ times the Dirichlet GFF. Later Berestycki and Liu showed [8, Theorem 5.1] that the mean of $2\lambda$ times the height function converges to a harmonic function with piecewise constant boundary conditions with alternating jumps of size $\pm\lambda$ located at the points of the Temperleyan corners (in our case these are the two extremities of $\partial U\cap(\{0\}\times\mathbb{R})$) plus a term that involves the winding of the boundary of $U$. Since $U$ and $\bar{U}$ have opposite windings of the boundary, they cancel out when the two height functions are added, and we can conclude that the mean of $2\lambda$ times the height function of the superimposed configuration converges to the harmonic function in $U$ with $\lambda$ boundary conditions on $\partial U\cap(\{0\}\times\mathbb{R})$ and zero elsewhere. Since the sum of two independent fields that are both $1/\sqrt{2}$ times the Dirichlet GFF (coming from the upper and lower domains) is ($1$ times) the Dirichlet GFF, we can conclude that $2\lambda$ times the height function of our model converges to the corresponding field described in [18], i.e. a GFF in $U$ with boundary conditions equal to $\lambda$ on $\partial U\cap(\{0\}\times\mathbb{R})$ and zero on the rest of $\partial U$.
@@
We conjecture that the arcs in both models converge to the ALE in $U$ (emanating from $\partial U\cap(\{0\}\times\mathbb{R})$ as in [18]). This extends the conjecture of [7] which is the same statement only for the first “folded” model. Our conjecture is strongly evidenced by the discussion above. However, convergence of the height function is known not to be enough to conclude convergence of interfaces in the double dimer model [16, 12, 5]. Finally let us mention that the two models yield different measures in the discrete setting, when one forgets the orientation of loops and arcs. Indeed, in the first model each nontrivial arc (that is not a single edge) comes with a combinatorial factor of two corresponding to the two different orientations, and it comes with a factor of one in the second model. The number of nontrivial arcs is not constant (unlike the number of all arcs) and hence the measures are different. Nonetheless we conjecture that the two collections of arcs have the same scaling limit.
@@
The main observables of interest in both models will be $n_{\varepsilon}(z)$ – the number of arcs that surround a point $z$, i.e. arcs that separate $z$ from the part of the boundary with no arcs $\partial U\setminus(\{0\}\times\mathbb{R})$, and $o_{\varepsilon}(z)$ – the parity of $n_{\varepsilon}(z)$. Our main result can be summarised as follows.
@@
One of the main conceptual difficulties in approaching the problem of studying the arcs themselves is the issue of separating the loops from the arcs in the counting arguments. This is achieved by developing a novel variant of Kenyon’s formula for double dimers with arcs, and afterwards choosing the right weights (or more precisely SL(2) connections) in this formula. Afterwards an involved analysis of the scaling limit of the formula by means of discrete complex analysis techniques is required.
We note that the works [16, 12, 5] study multi-point functions of topological observables in the double dimer model. Basok and Chelkak [5] showed that these observables contain enough information to identify the law of the loops. In our case, the fact that we want to disregard all the loops from the picture restricts greatly the number of useful topological observables, and we can only treat the one-point (or two-point) function. However, it will be very interesting to develop the methods of [16, 12, 5] to study the full collection of both arcs and loops, that (based on the conjecture of Kenyon [16]) and our conjectures should converge to the ALE with an independent CLE4 in its complement.
Below we include a table summarising the main properties of our two dimer models.
Folded dimer model Shifted dimer model
Measure notation $\mu^{1}_{\varepsilon}$ $\mu^{2}_{\varepsilon}$
Height function b.c. in discrete Varies Alternates between 0 and 1
Limit GFF b.c. on bdry arc Neumann Dirichlet
Orientation of arcs in discrete i.i.d. Bernoulli Deterministic
Weight on arcs in discrete $2^{\#arcs}$ $1^{\#arcs}$
Conjectured limit of arcs ALE ALE
@@ Acknowledgements
The authors extend their warm thanks to Avelio Sepúlveda for his contribution of Appendix A. ML and LR thank Misha Basok for useful discussions during the program Geometry, Statistical Mechanics, and Integrability at IPAM, UCLA. ML and KR were supported by the FWF Standalone grant Spins, Loops and Fields P 36298 and the FWF SFB Discrete Random Structures 10.55776/F1002. KR was supported by the Academy of Finland Centre of Excellence Programme grant number 346315 “Finnish centre of excellence in Randomness and STructures” (FiRST). LR was partially supported by the DIMERS project ANR-18-CE40-0033 funded by the French National Research Agency.
@@ 2.1. Setup and formula
We begin with some definitions. Let $G=(V,E)$ be a finite, connected, bipartite, planar graph. Its vertices $V$ are partitioned into white and black vertices. Let $\partial$ be a strict subset of the vertices incident to the external face of $G$, which contains the same number of black and white vertices. We form a graph $G^{\times}$ which can be described as taking two copies of $G$, identifying the vertices in the two copies on $\partial$ (and not the rest of the graph), and then adding “diagonal” edges which join a vertex $u$ in one copy of $G$ to a vertex $v$ in the other copy, whenever $u$ and $v$ form an edge in $G$. One can think of the two copies of $G$ being “folded” together along $\partial$. See Figure 2 for a small example.
Figure 2. A very simple example of the graph $G^{\times}$ (right), given a graph $G$ (left). The set $\partial$ is in this case the single red vertex.
@@
A dimer configuration on $G^{\times}$ is a set of edges $m\subset E^{\times}$ such that every vertex in $V^{\times}$ is incident to exactly one edge in $m$. The projection of a dimer configuration $m$ is the set of edges $p(m)\subset E$ obtained by projecting all edges of $m$. In the case $\partial=\varnothing$, the projection of a dimer configuration on $G^{\times}$ is a configuration of disjoint closed loops and double edges on $G$: it is a double dimer configuration. In the case $\partial\neq\varnothing$, the projection of a dimer configuration on $G^{\times}$ is a configuration of disjoint closed loops, double edges, and also paths from $\partial$ to $\partial$, which we call arcs (the arcs can be of length 1). Let $\Omega$ denote the set of such configurations of loops, double edges, and arcs on $G$.
Let $\Phi$ be some $\mathrm{SL}_{2}(\mathbb{C})$ connection in the bulk of $G$. That is, $\Phi$ assigns each directed edge uvuv ($u,v\in V\setminus\partial$) a matrix $\phi_{u,v}\in\mathrm{SL}_{2}(\mathbb{C})$, such that $\phi_{v,u}=\phi_{u,v}^{-1}$. For a loop $C=\{e_{1},\dots,e_{2n}\}$ in $G$, define the monodromy of $\Phi$ around $C$, $\phi_{C}$ to be the product $\phi_{e_{1}}\circ\cdots\circ\phi_{e_{2n}}$. Note this depends on the choice of $e_{1}$ and the orientation or the loop; $\mathrm{tr}(\phi_{C})$, however, does not. For an edge $\{u,w\}$ with $w\in\partial$ and $u\in V\setminus\partial$, let $\psi_{\{u,w\}}\in\mathbb{C}^{2}$. Fix some basis of each copy of $\mathbb{C}^{2}$ so that each $\phi_{v,u}$ can be written as a matrix, and each $\psi_{\{u,w\}}$ as a column vector.
@@
It is well known that such phases exist (see for example Theorem 4.1 of [10]) and in fact can be taken to be real i.e. signs $\pm 1$ on the edges. One can show that (1) holds for all loops surrounding an even number of vertices. Moreover if one replaces $\xi_{x,y}$ by $\varepsilon(x)\varepsilon(y)\xi_{x,y}$ for some function $\varepsilon:V\to\{\pm 1\}$, the new phases are also a set of Kasteleyn phases: this is called a gauge transform. Let $\bar{\partial}$ be a connected, strict subset of the boundary vertices, with $\partial\subset\bar{\partial}$. By using a guage transform, one can assume that the phases on the boundary edges (incident to the external face) alternate between +1+1 and 1-1 in $\bar{\partial}$. We fix such real phases $\xi$ for the rest of this section.
@@
We assign an ordering to $V^{\times}$ which will be used in the Pfaffian of Proposition 2.1. Let the ordering be: first the boundary vertices, ordered as they appear clockwise in $\overline{\partial}$, then the white vertices in some order such that each $(u,1)$ is directly followed by $(u,2)$, and then the black vertices, again in some order such that $(v,1)$ is directly followed by $(v,2)$. We can write our assumption that the Kasteleyn phases alternate on the boundary as $\xi_{w,b}=(-1)^{\mathds{1}\{w>b\}}$ for all edges linking two vertices of $\overline{\partial}$.
@@
Figure 3. An example of a graph $G$, where we take $\partial=\overline{\partial}$ to be the edges on the $x$-axis, and real Kasteleyn weights $\xi$ which alternate in sign along $\overline{\partial}$ as $\xi_{w,b}=(-1)^{\mathds{1}\{w>b\}}$, where the ordering on $\partial$ increases clockwise. (We only write the 1-1 phases in the diagram for clarity; the rest are 1).
@@
(thinking of $b,w\in\mathbb{C}$), that is, at each white vertex, the weights of the four adjacent edges starting at the right-going edge and proceeding anticlockwise are 1,i,1,i1,i,-1,-i respectively. Using the gauge function $\varepsilon(u)=(-i)^{\mathds{1}\{u\text{ has odd vertical coordinate}\}}(-1)^{\mathds{1}\{u\ \text{is white}\}}$, and the real Kasteleyn phases $\xi$ depicted in Figure 3, $\zeta_{u,v}:=\varepsilon(u)\varepsilon(v)\xi_{u,v}$ are the Kasteleyn phases of discrete holomorphy of Equation (3). Remark 2.3 gives us the corollary of Proposition 2.1 which we will use in the following sections.